OpenStudy (anonymous):

Can you simplify an expression with a 0 exponent

OpenStudy (ipwnbunnies):

Yes. \[a^{0} = 1\]

OpenStudy (ipwnbunnies):

Where a can be a number, a variable, a giraffe even.

OpenStudy (anonymous):

WHY WOULD YOU PWN A BUNNY ;"(

OpenStudy (anonymous):

yes, because a^o= 1

OpenStudy (zarkon):

what is \(0^0\)?

OpenStudy (anonymous):

1. There's a lot of evidence for it, but some people seem to think it implies that 0/0 = 1.

OpenStudy (zarkon):

it is undefined

OpenStudy (zarkon):

there are valid reasons to want it to be 1...but there are also valid reasons to want it to be 0

OpenStudy (anonymous):

Proven doesn't mean true. Here is why I think it is 1. This is one of the many sources I pull my opinion from. http://www.math.hmc.edu/funfacts/ffiles/10005.3-5.shtml Also, computer science finds it convenient that 0^0 = 1.